OpenStudy (anonymous):

help The height of an equilateral triangular prism increases by one unit. The new lateral area is more than the original by how much? the area of the base the height the perimeter of the base the area of one lateral face

OpenStudy (anonymous):

my answer is c

OpenStudy (cwrw238):

lateral area = perimeter * height of prism = s * h new area = s (h + 1) = sh + s answer is perimeter of base

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks you so much

OpenStudy (cwrw238):

yw